OpenStudy (anonymous):

Four weights of 1, 3, 9, and 27 pounds sit on a table along with two identical boxes. One of the boxes contains an object that weighs X pounds, where X is an integer between 0 and 40. Explain how, for any such X, it is possible to place some or all of the weights in the boxes so the contents of the boxes then have the same weight.

OpenStudy (zarkon):

I'll give an example and see if you can get it from there if X=38=1*27+1*9+2*1 =27+9+3-1 38+1=27+9+3 you can do this for every number between 0 and 40